# Transportability of Prognostic Markers: Rethinking Common Practices through a Sufficient-Component-Cause Perspective

### v2

@@ Abstract
We modeled population-generating scenarios similar to those presented in Figure 1. Because in many realistic settings all three causes are likely to vary, we explored this setup with more depth. We modeled scenarios where changes in the three causes had different strengths of correlation with each other (ranging from changes being uncorrelated to fully correlated). We also modeled a ‘maximum entropy’ scenario where the three causes have independent uniform distributions; albeit unrealistic, this scenario represents maximum theoretical randomness across populations (therefore, marker performance is entirely untransportable)27. For each scenario, we simulated random pairs of source and target populations according to a given population-generating mechanism. When a cause was made variable, it was assumed that the logit-transformed probability of that cause had a standard Normal distribution (e.g., $\mbox{logit}(P_{T})\sim\mbox{Normal}(0,1)$ when where $P_{T}$ was variable across populations). For each pair, we applied the following three transportation strategies: by accuracy (SE and SP), by predictive values (PPV and NPV), and via the proportional-odds assumption explained previously. Divergence metrics were averaged across 10,000 simulations for each scenario.
@@
Results are provided in Figure 2. This confirms that if the outcome prevalence variation is entirely due to $T$, PPV and NPV are fully transportable ($D_{KL}=0$). However, under other scenarios, transportation by predictive values resulted in information loss, sometimes substantially. Transportation by accuracy reduced information loss in some scenarios, but increased in others. This indicates that prevalence adjustment is not guaranteed to improve the transportability of a prognostic marker. In none of the scenarios was the loss zero, indicating that none of the modeled scenarios were compatible with stable SE and SP - again, due to the complex requirements for the stability of these metrics. The proportional-odds assumption also reduced information loss in some scenarios, but increased it in others. For the scenarios where the three causes changed together, transportation by accuracy generally reduced information loss compared with transportation by predictive values. In such scenarios, transportation via the proportional-odds assumption performed better than both conventional transportability methods.
@@ Discussion
We constructed a simple causal model under the SCC framework to study the transportability of prognostic markers that are partially informative about an outcome of interest but also generate false-positive and false-negative results. Several observations from this theoretical exploration deserve being highlighted. First, the common practice of transporting the marker by predictive values is questionable, as it relies on the strict assumption that the prevalences of causes other than those represented by the marker remain stable across populations. Further, we showed that the conventional prevalence adjustment method is the exact implementation of Bayes’ rule, thus changing the locus of transportability from predictive values to accuracy metrics (sensitivity and specificity). Nevertheless, no easily explainable or biologically plausible mechanism is likely to generate fully transportable accuracy metrics. In our explored scenarios regarding the distribution of causes across populations, transporting a prognostic marker by accuracy was not universally better than transporting by predictive values. However, in scenarios where causes were positively correlated, which is likely to be common, transporting by accuracy did reduce information loss compared with transportability by predictive values. However, in the same scenarios, the SSC-based ‘proportional odds’ method of transportation performed better.
@@
Our simulations were related to a selected scenarios and particular ways the causes vary across populations. Still, they should be sufficient to question some common assumptions and practices. These include the practice of advertising prognostic risk questions without emphasizing their dependence on prevalence, or considering sensitivity and specificity as intrinsic properties of markers. While these insights are from studying binary markers, the core findings can be extended to continuous markers and multi-variable risk equations. Transforming such markers ‘as is’ is equal to assuming that stratum-specific predictive values ($P(D|T)$) remain constant. This is equal to attributing variations in prevalence entirely to variations in $P_{T}$ (aka case-mix). Consider a multi-variable risk score such as the QRISK for cardiovascular diseases28. Even though this model includes up to 22 predictors, is it plausible that other causal pathways to cardiovascular diseases that might vary across populations, including environmental exposures, subtle lifestyle choices, access to and quality of preventive care, and yet-to-be-discovered genetic risk factors, are the same across populations? This assumption is needed for claiming that QRISK predictions, developed using primary care UK data, are transportable to other settings. On the other hand, conventional prevalence adjustment would indicate that stratum-specific likelihood ratio, i.e., $P(x|D=1)/P(x|D=0)$ would be transportable. This assumption will hold if the distribution of predicted risks among those who will experience the outcome and those who will not remains the same across diverse populations. This condition places very specific constraints on how the causes of diseases should vary across populations.
@@
We focused on prognostic markers given the recent debates on transportability of prognostic information and the merit of prevalence-adjustment to improve transportability. The application of SCC framework for diagnostic markers is also important and deserves its own airing. For such markers, the path of causality if from the disease to the test. As such, an equivalent setup to our reference setup for a diagnostic marker would be $T=(D\lor V)\land U$. The conventional method of transporting diagnostic markers is by accuracy (e.g., the Bayes’ rule, which combines accuracy metrics from the source population with disease prevalence from the new population). From this setup, it is immediately obvious that transportation by accuracy requires that universal and alternative causes $U$ and $V$ remain stable across populations. This assumption is similar to the assumption of transportability by predictive values for prognostic markers. This strong requirement can explain why sensitivity and specificity tend to vary across populations29–31. How the SCC framework can inform transportability of diagnostic markers needs to be pursued in future work.
@@
How can these findings inform practice? If reliable, context-specific information is available on the distribution of specific predictors and their relationships with the outcome and with each other, methods based on causal graphs can be used to build transportable models or to design tailored transportation strategies11. In contrast, when only general information, such as outcome prevalence or test positivity rates, are available, the SCC framework can be used to formulate transportability algorithms that utilize such broad information. We question the contemporary practice of advertising risk prediction models for outcome risk as transportable. We believe risk equations should explicitly mention the outcome risk in their source population, and adjustment for prevalence should be advocated as a necessary step before using the equation in a population with different outcome prevalence. Yet, while conventional prevalence adjustment may serve as a reasonable choice, it cannot be assumed to consistently improve performance. Instead of fixed transportability rules, the SCC framework offers a foundation for updating marker performance in ways that are more grounded in mechanistic interactions among potential causes. This results in new transportability algorithms based on assumptions on how underlying causes vary, rather than keeping performance metrics, which are emergent properties of such causes, constant. An example is the proportional-odds prevalence adjustment method discussed above. It is motivated by the notion that oftentimes differences in outcome prevalence are multi-factorial. An agnostic stance that treats all causes as equally responsible is more defendable than blaming only one (as in transportation by predictive values) or assuming they move in unexplainable ways just to keep sensitivity and specificity constant (as in transportability by accuracy). Interestingly, in our numerical experiments, whenever conventional prevalence-adjustment performed better than transportability by predictive values, the SCC-based proportional-odds method performed better than both, indicating it might be a better choice. Of course, variations in this method are also possible. Imagine, for example, that in addition to knowing target prevalence, we have knowledge of marker positivity rate ($P_{T}$) in the new population (e.g., from an independent cross-sectional study). One can then directly adjust $P_{T}$ to match this value, and then apply a common odds-ratio to other causes to match target prevalence.
@@
We conclude this paper by providing a few areas for further inquiry. Whether SCC-informed transportation algorithms such as the proportional-odds method actually improve transportability can be tested in empirical studies. Our arguments were based on precise knowledge of outcome prevalence in the target population. To what extent using a noisy estimate of prevalence will help or harm transportability needs to be investigated. The issue of uncertain prevalence estimate is also applicable to conventional method. Empirical studies on conventional prevalence adjustment also estimated the prevalence from the sample that was subsequently used to assess marker performance - effectively taking its value as known2,3. Further, our explorations were for when performance metrics are derived from a single source population. Without knowing how populations differ from each other, the choice of transportation method will require assumptions on the distribution of causes. On the other hand, when contingency tables from multiple populations are at hand, this choice can be learned from the data. This results in SCC-based model specifications for meta-analysis of marker performance studies, where the estimands are the parameters that govern the joint distribution of $T$, $U$, and $V$. This will provide an alternative to modeling the joint distribution of prevalence, sensitivity, and specificity (or prevalence and predictive values)32,33. The more explanatory nature of SCC-based specification might provide better fit to the data. In addition to pooled estimates of performance metrics and their predictive distribution for a new population, this approach provides an overall estimate of the degree by which universal and alternative causes are responsible for between-population variations, which might be of secondary interest.
@@
  1. 1.
    Map from {TP,FP,FN,TNTP,FP,FN,TN} to $P_{T},P_{U},P_{V}$} (see text)
  2. 2.
    Find the odds-ratio $x$ such that when applied to $P_{T}$, $P_{U}$, $P_{V}$, results in updated probabilities $P_{T}^{*}$, $P_{U}^{*}$, $P_{V}^{*}$ that correspond to the desired prevalence. Note that prevalence is equal to $P_{U}^{*}(P_{T}^{*}+P_{V}^{*}-P_{T}^{*}P_{V}^{*})$.
Of note: $x$ can be found as the real root of the cubic equation $ax^{3}+bx^{2}+cx+d=0$ with
@@
which can be solved using standard methods.

### v3

@@ Abstract
We modeled population-generating scenarios similar to those presented in Figure 1. Because in many realistic settings all three causes are likely to vary, we explored this setup with more depth, covering scenarios were causes had positive correlation of varying strengths. We also modeled a ‘maximum entropy’ scenario where the three causes have independent uniform distributions; albeit unrealistic, this scenario represents maximum theoretical randomness, and thus places an upper limit on information loss30. For each scenario, we simulated 10,000 random pairs of source and target populations. For each pair, we applied the following transportation strategies: by accuracy, by predictive values, and via the proportional-odds assumption explained previously.
@@
Results are provided in Figure 2. Overall, they show that no transportability method is universally better than others. They confirm that if the outcome prevalence variation is entirely due to $T$, PPV and NPV are fully transportable ($D_{KL}=0$). However, under other scenarios, transportation by predictive values resulted in information loss, sometimes substantially. For transportability by accuracy, in none of the scenarios was the loss zero, indicating that none of the modeled scenarios were compatible with stable SE and SP - again, due to the complex requirements for the stability of these metrics. Notably, for the scenarios where the three causes changed together, the new, proportional-odds assumption performed better than both conventional transportability methods.
@@ Discussion
We constructed a parsimonious causal framework to study the transportability of prognostic markers. Several observations from this theoretical exploration deserve reiterating. First, the common practice of transporting prognostic markers by predictive values relies on the strong assumption that the prevalences of universal and alternative causes remain stable across populations. We showed that the conventional prevalence-adjustment method is the exact implementation of Bayes’ rule, thus changing the locus of transportability to accuracy (sensitivity and specificity). Nevertheless, no easily explainable or biologically plausible mechanism is likely to generate fully transportable accuracy metrics. In our explored population-generating scenarios, transporting a prognostic marker by accuracy was not universally better than transporting by predictive values. However, in scenarios where causes were positively correlated, which is likely to be common, transporting by accuracy reduced information loss compared with transportability by predictive values. However, in the same scenarios, the new ‘proportional odds’ method of transportation performed better.
@@
Our numerical results were related to a selected ways the causes vary across populations. Still, they should be sufficient to question some common assumptions and practices, including the practice of advertising prognostic risk questions without emphasizing their dependence on case-mix and outcome prevalence, and considering sensitivity and specificity as intrinsic properties of markers. While these insights are from studying binary markers, the core findings can be extended to continuous markers and multi-variable risk equations. Transforming such markers ‘as is’ is equal to assuming that stratum-specific predictive values ($P(D|T)$) remain constant. This is equal to attributing variations in prevalence entirely to variations in $P_{T}$. Consider a multi-variable risk score such as the QRISK3 for cardiovascular diseases31. Even though this model includes up to 22 predictors, is it plausible that other causes of cardiovascular diseases, including environmental exposures, lifestyle choices, access to and quality of preventive care, and genetic risk factors, are the same across populations? This assumption is needed for claiming that QRISK3 predictions, developed using primary care UK data, are transportable to other settings. On the other hand, conventional prevalence-adjustment would indicate that stratum-specific likelihood ratios would be transportable (i.e., $P(T=x|D=1)/P(T=x|D=0)=c$ for all $x$). This assumption will hold if the distributions of predicted risks within outcome status strata remain the same. This condition places very specific constraints on how the causes of diseases should vary across populations.
@@
We focused on prognostic markers given the recent debates on transportability of prognostic information and the merit of prevalence-adjustment to improve transportability. The application of this framework for diagnostic markers is also important and deserves its own airing. For such markers, the path of causality if from the disease to the test. As such, an equivalent setup to our reference setup for a diagnostic marker would be $T=(D\lor V)\land U$. From this setup, it is immediately obvious that transportation by accuracy (the default approach for diagnostic markers) requires that universal and alternative causes ($U$ and $V$) remain stable across populations. This strong requirement can explain why sensitivity and specificity tend to vary across populations2–4. How the SCC framework can inform transportability of diagnostic markers needs to be pursued in future work.
@@
How can these findings inform practice? we question the contemporary practice of advertising risk prediction models for outcome risk as transportable. If reliable, context-specific information is available on the distribution of predictors and their relationships with the outcome and with each other, methods based on causal graphs can be used to build transportable models or to design tailored transportation strategies13. In contrast, when only general information, such as outcome prevalence or test positivity rates, are available, the SCC framework can be used to formulate transportability algorithms that utilize such broad information. Instead of fixed transportability rules, this framework offers a foundation for algorithms that reflect our assumptions on how underlying causes vary, rather than keeping performance metrics, which are emergent properties of such causes, constant.
@@
We conclude this paper by providing a few areas for further inquiry. Whether SCC-informed transportation algorithms such as the proportional-odds method actually improve transportability can be tested in empirical studies. Our arguments were based on precise knowledge of outcome prevalence in the target population. To what extent using a noisy estimate of prevalence will help or harm transportability needs to be investigated. The issue of uncertain prevalence estimate is also applicable to conventional method. Empirical studies on conventional prevalence adjustment also estimated the prevalence from the sample that was subsequently used to assess marker performance - effectively taking its value as known5,32. Further, our explorations were for when performance metrics are derived from a single source population. Without knowing how populations differ from each other, the choice of transportation method will require assumptions on the distribution of causes. On the other hand, when contingency tables from multiple populations are at hand, this choice can be learned from the data. This results in SCC-based model specifications for meta-analysis of marker performance studies, where the estimands are the parameters that govern the joint distribution of $T$, $U$, and $V$. This will provide an alternative to modeling the joint distribution of prevalence, sensitivity, and specificity (or prevalence and predictive values)33,34. The more explanatory nature of SCC-based specification might provide better fit to the data. In addition to pooled estimates of performance metrics and their predictive distribution for a new population, this approach provides an overall estimate of the degree by which universal and alternative causes are responsible for between-population variations, which might be of secondary interest.
@@
Step 1. Map from {TP,FP,FN,TNTP,FP,FN,TN} to $P_{T},P_{U},P_{V}$} (see text)
Step 2. Find the odds-ratio $x$ such that when applied to $P_{T}$, $P_{U}$, $P_{V}$, results in updated probabilities $P_{T}^{*}$, $P_{U}^{*}$, $P_{V}^{*}$ that correspond to the desired prevalence. Note that prevalence is equal to $P_{U}^{*}(P_{T}^{*}+P_{V}^{*}-P_{T}^{*}P_{V}^{*})$.
Given that applying an odds-ratio $x$ to probability $p$ can be written as as the function $\frac{px}{1-p+px}$, we have
$P_{T}^{*}(x)=\frac{P_{T}^{*}x}{1-P_{T}^{*}+P_{T}^{*}x}$ (similar for $P_{U}$ and $P_{V}$).
This, we should solve for $x$ in $f(x)=\pi^{*}$ where
$f(x)=P_{U}^{*}(x)[P_{T}^{*}(x)+P_{V}^{*}(x)-P_{T}^{*}(x)P_{V}^{*}(x)].$
We consider non-degenerate scenarios where $0<P_{T}<1$, $0<P_{U}<1$, and $0<P_{V}<1$. Given that $f(x)$ is continuous, $f(0)=0$, and $\lim_{x\to\infty}f(x)=1$, there is always at least one real solution for $x$. Further, given that
  • $\frac{dP_{T}^{*}(x)}{dx}>0$, $\frac{dP_{U}^{*}(x)}{dx}>0$, and $\frac{dP_{V}^{*}(x)}{dx}>0$, and
  • $\frac{\partial f(x)}{\partial P_{T}^{*}(x)}>0$, $\frac{\partial f(x)}{\partial P_{U}^{*}(x)}>0$, $\frac{\partial f(x)}{\partial P_{V}^{*}(x)}>0$,
$\frac{df(x)}{dx}>0$ for all $x$. As such, this solution is unique. This solution can be found using univariate root finding methods (e.g., uniroot() in R).
Of note, re-arranging the terms reveals that $x$ can also be expressed as the real root of the cubic equation $ax^{3}+bx^{2}+cx+d=0$ with
@@
which can be solved using standard methods (e.g., polyroot() in R).